Circular No. 5890

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

HARD X-RAY TRANSIENT B. A. Harmon, S. N. Zhang, G. J. Fishman, and C. A. Wilson, Mar-

shall Space Flight Center, NASA, report for the COMPTON Observatory BATSE Team: “An increase in hard x-ray flux in the band 20-230 keV has been observed since Nov. 10 in the Aquila-Scutum region using Earthoccultation monitoring. Maximum intensity from preliminary data was on Nov. 13 at ∼ 100-150 mCrab, with a hard spectrum extending above 200 keV. Its location is not well-determined at this time; the error box is ± 0^{o}.5 degrees along a line between the locations α = 18^{h}28^{m}, δ = +8^{o} and 18^{h}52^{m}, −4^{o}. Assuming that the source remains bright for several more days, the location accuracy should improve.”

X-RAY NOVA IN VELA M. Della Valle, European Southern Observatory; and S. Benetti, De-

partment of Astronomy, University of Texas at Austin, report: “We have found the probable optical counterpart of the x-ray source in Vela (cf. IAUC 5864, 5886) at V = 14.6, B–V = +0.3, V–R = +0.1, near the edge of the ASCA error box (IAUC 5888), on a CCD frame obtained on Nov. 17.3 UT with the ESO/MPI 2.2-m telescope (+ EFOSC2) at La Silla. A preliminary position measured on the CCD frames is α = 10^{h}11^{m}32^{s}, δ = −44^{o}49′41′′ (equinox 1950.0). Analysis of a spectrogram (range 470-690 nm, resolution ∼ 0.7 nm) secured with the same equipment shows a relatively blue continuum, with a broad Hα emission line (FWHM ∼ 3 nm) clearly separated into two components measured at 655.6 and 657.3 nm. The emission line due to Hβ is also visible. Inspection of the ESO B Schmidt plate shows the probable progenitor at B ∼ 19.”

COMET MUELLER (1993a) Total visual magnitude estimates: Sept. 18.23 UT, 10.3 (C. S. Morris,

Pine Mountain Club, CA, 0.26-m reflector); Oct. 10.19, 9.3 (G. W. Kronk, Troy, IL, 20×80 binoculars); 25.38, 9.1 (J. E. Bortle, Stormville, NY, 0.32-m reflector); Nov. 4.75, 9.1 (A. Diepvens, Balen, Belgium, 0.15-m refractor); 16.11, 9.2 (C. E. Spratt, Victoria, BC, 0.20-m reflector).

1993 November 17 Daniel W. E. Green